OpenStudy (anonymous):

Which ratio is equivalent to 7 : 9? nine over seven 14 to 18 21 : 3 7 to 27

OpenStudy (calculusxy):

It is 14:18 becuase 7*2=14 and 9*2=18

OpenStudy (debbieg):

You are looking for \[\dfrac{ 7 }{ 9 }=\dfrac{ ? }{ ? }\] For two ratios to be equivalent, both the numerator and denominator have to be multiplied by the SAME number: \(\dfrac{ 7 }{ 9 }=\dfrac{ 7a }{9a }\) So you are looking for a num'r that is the same multiple of 7 that the den'r is of 9.

OpenStudy (calculusxy):

Also known as 14 to 18.

OpenStudy (debbieg):

No, I think "nine over seven" is the first of the 4 multiple choice answers, @calculusxy

OpenStudy (calculusxy):

Oh okay.

OpenStudy (calculusxy):

Did you find your answer?